OpenStudy (deoxna):

A, B, C are angles of a triangle. Show that tan A + tan B + tan C = A tan B tan C

OpenStudy (deoxna):

This is exactly how the question is written. However, I suspect that the A must be actually be (tan A) since I don't know of anyway you can "extract" an angle from a tangent (other than arctan, but that would entail additional variables). Either way, I've been trying for a while and can't find an answer,

OpenStudy (chihiroasleaf):

yes.., I think it's should be tan A + tan B + tan C = (tan A)(tan B) (tan C) since A, B and C are angles of a triangle, then we have, A+B+C=180 A+B = 180-C now, tan(A+B) = tan(180-C)= -tan C now, use the identity formula of tan(A+B) , then simplify

OpenStudy (deoxna):

Thanks alot!

OpenStudy (chihiroasleaf):

have you got it? :)

OpenStudy (deoxna):

yep, thanks for all the help